How many ways are there to answer 15 multiple choice questions with 4 choices each?
There are 4*4*4... 15 times
so ans would be = 415
How many ace of diamond in a deck of 52 cards?
There is only one ace of diamonds in a normal deck of 52 cards.
1 out of 6
* * * * *
Total rubbish.
There are 11 possible sums - the numbers 2 to 12. So if you throw the dice 12 times, the first 11 can be different but the 12th must be a repeat.
What is the probability of rolling the sum of 10 with 2 dice?
Well there is 36 different possibilities with rolling 2 6 sided dice.
The probability of rolling the sum of 10 with 2 die is 4/36 or 1/8 chance.
What is probability of drawing a king and a queen from a deck of cards?
The probability of drawing a king and a queen from a standard 52 card deck is 32 in 2652, or 8 in 663, or about 0.012066.
The first king or queen is 8 in 52. The second card is 4 in 51. Simply multiply 8/52 by 4/51.
What event has a probability of 0?
-- The freezing over of Hell.
-- A confirmed sighting of flying pigs.
etc.
Besides anything that is impossible, as noted above, the simultaneous occurrence of two mutually exclusive events has a probability of zero: It is raining and it is not raining, it is on and it is off, etc.
How do you find ratio 8 and 7?
8/7=1.1428571428571428571428571428571
so the ratio of 8:7 means it takes 1.1428571428571428571428571428571 sevens to equal 8
Why is it important to carry out a large number of trials in determining probability?
If you only carry out a few trials, then how can you know how many times a particular situation will occur? One has to do a lot of trials in order to determine how many times that situation will happen so he can conclude the probability he's looking for.
How many different three-letter initials can people have?
first space filled with A-z this implies that n(first)=26.
second space filled with all letters from a-z this implies that n(second)=26.
third space filled with all letters from a-z this implies that n(third)=26.
therefore by principle of counting
26x26x26=17576
What is the probability of two prime numbers when rolling two dice?
The probability of rolling two prime numbers on a standard pair of dice is 1 in 4, or 0.25. Take the probability of rolling a prime on one die, 3 in 6, or 1 in 2, or 0.5, and square it.
What are kissing kings in cards?
The kissing kings are the two that meet in the middle of a newly opened deck of cards: King of clubs and King of diamonds.
What is the probability of picking a king out of a pack of playing cards?
There are 4 kings in a pack of 52 cards so the odds are 1:13 (4/52).
What is the probability of rolling an odd number or an even number on a cube?
To answer this problem you first see how many even and odd numbers there are. There are 3 odd and 3 even. so the fraction for even and odd is 3/6 so then in simplified form it is 1/2. If they ask to express the probability as a decimal number it would be .5 or.50
It all will depend on what kind of (recessive or dominant) alleles are responsible for the colorblind characteristic and what kind of alleles do the parental genes have.
P(A|B)= P(A n B) / P(B)
P(A n B) = probability of both A and B happening
to check for independence you see if P(A|B) = P(B)
What is the chances of Rolling a 3 or 4 from two dice?
Assuming you are rolling a six-sided dice, it is 1/3 for one dice and 2/3 for two dice.
How many outfits can you make with 4 sweaters 6 dress shirts and 3 pairs pants?
There are 30 outfit possibilities. Normally for a problem like this you just have to multiply all the numbers, but the trick to this question is to first add the number of shirts. If you don't, then some of your outfits will be a sweater and dress shirt with no pants, which isn't an outfit. So you have a total of 10 shirts and 3 pairs of pants, and when you multiply them you get a total of 30 outfits.
What is High Impact and Low Probability Analysis?
It's an analytic technique similar to "what if" analysis, that looks at events which are regarded as unlikely, however; would have a significant impact if they were too happen. September 11 is an example of an event that was considered unlikely pre 9/11, but has obviously had a huge impact. High Impact/Low Probability analysis seeks to identify the possibility of events like this before they happen in an effort to prevent 9/11-like events.
What is the probability of rolling at least one six with one roll of a pair of dice?
To find the probability of rolling at least one six with one roll of a pair of dice, we need to know two things. First we need to know how many different possible outcomes there are when we roll the dice one time. We know that there are 36 different possible outcomes. They are 1-1, 1-2, 1-3,..., 6-4, 6-5 and 6-6.
We also need to know how many different times at least one six can appear. These will include, 1-6, 2-6, 3-6, 4-6, 5-6, 6-6, 6-1, 6-2, 6-3, 6-4, and 6-5, for a total of 11 possible outcomes with at least one six. (Note that the 6-6 possible outcome is included in the list because it fits the requirement of "at least one six" with a given single pass of the dice.
The odds of rolling at least one six in one pass with a pair of dice are 11 in 36. To find the probability of this, we need to divide the 11 by the 36, and this gives us a probability of 11/36 or 0.3055555....
Remember that probability is a ratio, and it's the ratio of the number of possible times a given outcome can appear to the total number of possible outcomes that could appear. This ratio will range from zero to one, inclusive. A probability of zero means that a given outcome cannot occur. Like rolling a 13 with one pass of a regular pair of dice. That can't happen, so the probability of it happening is zero.
We also understand that if the probability of something happening is one, then that outcome must happen. Like the probability of rolling an odd or an even number. We know that an odd or an even number must appear, so the probability of that happening is one.
What is the probability of flipping a coin twice and getting two heads?
The probability is 25%.
The probability of flipping a coin once and getting heads is 50%. In your example, you get heads twice -- over the course of 2 flips. So there are two 50% probabilities that you need to combine to get the probability for getting two heads in two flips.
So turn 50% into a decimal --> 0.5
Multiply the two 50% probabilities together --> 0.5 x 0.5 = 0.25.
Therefore, 0.25 or 25% is the probability of flipping a coin twice and getting heads both times.